Reports 1-1 of 1 Clear search Modify search
Detector Characterisation (Glitches)
mwas - 20:20 Thursday 20 November 2025 (68209) Print this report
New very loud glitches - 15 minute glitches

While looking at lock acquisition I have noticed today a new family of glitches that have not been happening before Monday.

Figure 1 shows a few of those glitches today, they are extremely loud so they are visible directly in the B1 DC time series. They cause saturation of the B1 photodiodes and often kicks in the SDB1 alignment that are caused by these B1 photodiode saturations. The spacing looks very regular, with one glitch every 15 minutes.

Figure 2 these glitches happen in multiple locks, and might be causing lock losses.

Figure 3 looking at the timing of the lock losses, it appears roughly at the time of when the next 15 minute glitch should occur

Figure 4 these glitches have a shape of an oscilation at ~1.3Hz

These glitches will be a problem to keep the interferometer locked. I wonder if they are in addition to the 25 minute glitches, or if the 25 minute glitches have been replaced by a much louder and more frequent version.

Images attached to this report
Comments to this report:
direnzo - 1:27 Friday 21 November 2025 (68210) Print this report

I performed some quick analyses of these glitches and report a few preliminary observations below.

  • Although they appear very strong in B1, their SNR in the strain channel is about 220-270, roughly half that of the 25-minute glitches. Figure 1 shows the glitchgram of the strain data, where these glitches have a central frequency around 100 Hz.
  • Their waveform differs from that of the 25-minute glitches: they have a broader bandwidth and are shorter in duration, although they do show residual excitations after the peak: Figure 2.
  • They seem to have a preferred separation of roughly 14′45″ ± 2″ (to be confirmed with more statistics), but they can also occur very close to one another: Figure 3.
  • I ran UPV on both the strain and on SDB2_B1_PD1_D, but I did not find any evidence of correlations for these glitches (some insights for others). I will repeat the analyses over the next few days with larger statistics.
Images attached to this comment
mwas - 8:43 Friday 21 November 2025 (68212) Print this report

Figure 1. The problem can be already seen in CARM NULL 1F, in the trend data DARM has also glitches spaced by 15min before we reach LN3 (before there is power on B1)

Figure 2. Looking at raw data in CARM NULL 1F, the issue is actually more clearly visible in PRCL, in trend data the high frequency fluctuations of DARM are supressed, which is the reason in trend data it is more clearly visible in DARM.

Figure 3. In RDS 50Hz data the glitches can be clearly seen in PRCL, both in CARM NULL and in LN3

Figure 4. The problem was already present in the night from Nov 18 to 19, when we had the first longer CARM NULL lock after the electrical incident of Nov 17.

Figure 5. The problem was not presen in the morning of Nov 17 before the electrical incident.

A question is if the glitches remain synchronized to the ~15min spacing between different locks.

Another question if is there is any suspension glitches or environmental glitches that are coincident. By glitch, I mean a transient at ~1.3Hz, which means it will not be detected by omicron which typically looks only at higher frequencies.

 

Images attached to this comment
ruggi, boschi - 13:17 Friday 21 November 2025 (68213) Print this report

There are coincident glitches in the WE F7 control signals. They started on Monday, after the timing crisis. The involved signals are the LVDT's received from the coil driver board and reconstructed as error signals of the loops. The raw signals, sent to the master board from the LVDT board, look fine, while the recontructed signals are corrupted, so we guess that the issue is at the level of the coil driver, or in the communication between LVDT and CD.

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×